### v1

# Think Twice Before You Judge: Mixture of Dual Reasoning Experts for Multimodal Sarcasm Detection

Abstract.
Multimodal sarcasm detection has attracted growing interest due to the rise of multimedia posts on social media. Understanding sarcastic image-text posts often requires external contextual knowledge, such as cultural references or commonsense reasoning. However, existing models struggle to capture the deeper rationale behind sarcasm, relying mainly on shallow cues like image captions or object-attribute pairs from images. To address this, we propose MiDRE (Mixture of Dual Reasoning Experts), which integrates an internal reasoning expert for detecting incongruities within the image-text pair and an external reasoning expert that utilizes structured rationales generated via Chain-of-Thought prompting to a Large Vision-Language Model. An adaptive gating mechanism dynamically weighs the two experts, selecting the most relevant reasoning path. Experiments on two benchmark datasets show that MiDRE achieves superior performance over baselines. Various qualitative analyses highlight the crucial role of external rationales, revealing that even when they are occasionally noisy, they provide valuable cues that guide the model toward a better understanding of sarcasm.
Multimodal Sarcasm Detection, Mixture of Experts, Multimodal Semantic Understanding
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Multimedia information systemsccs: Computing methodologies Information extraction
Figure 1. Comparison between external knowledge used by existing approaches vs our approach.
1. Introduction
Sarcasm is a complex figure of speech where the intended meaning of an utterance often differs from the literal meaning. Due to its subtle and sophisticated nature, sarcasm is often employed as a tool for trolling or mocking others on social media platforms. Users intentionally exploit the gap between literal meaning and intended sentiment to convey irony, criticism, or humor in a veiled manner, making detection particularly challenging (Tindale and Gough, 1987; Averbeck, 2013; Maynard and Greenwood, 2014; Badlani et al., 2019; Ghosh et al., 2021). With the rise in multimedia content on these platforms, research towards the detection of image-text sarcasm has gained attention in recent years.
Understanding and analyzing sarcastic image-text posts poses a significant challenge, as their implicit meaning is not always conveyed directly through the surface-level text or visual content, frequently relying on external contextual information rather than just the content itself. For instance, in Figure 1(a), the caption claims a shortage of cabs, while the image clearly depicts an abundance of them. This form of sarcasm is direct and easily recognizable without requiring external knowledge. In Figure 1(b), the caption ”time to change the channel for some good news” mocks the idea of escaping a storm warning by simply switching channels. While still indirect, the sarcasm is relatively easy to grasp through description of visual content. However, in Figure 1(c), which suggests Halloween items might appear as early as March or April, relies on knowledge of seasonal retail patterns to detect the sarcasm, which makes it extremely difficult to detect.
In light of the above discussed examples, earlier works (Pan et al., 2020; Liang et al., 2021; Tian et al., 2023; Qin et al., 2023; Tang et al., 2024) rely solely on the image and text inputs without incorporating any external knowledge. While such models can effectively detect sarcasm in cases with explicit incongruities, such as in Figure 1(a), they often struggle when deeper contextual understanding is required. To address this, recent studies (Cai et al., 2019; Liang et al., 2022; Liu et al., 2022a) have introduced external knowledge in the form of image captions or object–attribute pairs. Although these cues can help in scenarios like Figure 1(b), they are typically shallow and generic, offering surface-level descriptions with limited relevance to the underlying sarcasm. As a result, these approaches fall short in more complex cases that require commonsense reasoning or background knowledge about social or cultural contexts, as illustrated in Figure 1(c).
Hence, we argue the need for rich contextual reasoning, where external knowledge is not only verbose but also structured into modality-specific reasoning, for gaining multiple viewpoints from the input. Unlike generic reasoning as used by earlier models, these reasonings should be sarcasm-oriented, offering clear and relevant cues. In designing our approach, we are guided by the following key observations:
  1. (1)
    Leveraging LVLMs for Rationales: Large Vision-Language Models (LVLMs), equipped with rich pretrained knowledge, can be prompted to generate explanatory rationales for determining whether an image-text pair is sarcastic. These rationales offer external contextual cues that go beyond the surface-level content of the input.
  2. (2)
    Multimodal dependence of sarcasm: Sarcasm is not only dependent on the contradiction between the image and text but can also be inferred from individual modalities independently. Therefore, it is crucial to incorporate rationales from the image modality, the text modality, and their combined interpretation.
  3. (3)
    Hallucination and bias in generated rationales: While LVLM-generated rationales can be informative, they may include hallucinations or factual errors. Relying on them blindly can mislead the model. Therefore, our approach should combine these external cues with cues from the input image-text for more reliable predictions.
To this end, we propose MiDRE (Mixture of Dual Reasoning Experts), a model designed to explicitly combine external reasoning cues with internal reasoning from input image-text for multimodal sarcasm detection. To realise observations (1) and (2), we generate external contextual knowledge from a pretrained Large Vision-Language Model (LVLM) using chain-of-thought (CoT) prompting, guiding it to produce step-by-step rationales: first from the image, then the text, and finally from the combined image-text input.
To alleviate the hallucination problem specified in observation (3), MiDRE introduces two complementary expert networks (internal expert and external expert) within an LLM backbone. The internal expert considers the raw image-text pair, capturing sarcasm from surface-level incongruities, without relying on external information. To capture sarcastic cues from external knowledge, external expert considers the image-text content as well as the rationales generated from the LVLM. A learnable gating module dynamically fuses the outputs from both experts, enabling MiDRE to adaptively balance between noisy external rationales and cues from the input image-text. This selective integration leads to more robust and interpretable sarcasm detection. Thus our key contributions are as follows:
  • We introduce MiDRE, a novel framework that uses external reasoning and the reasoning based on content, using dual expert networks, with an adaptive and interpretable gating mechanism for identifying sarcasm.
  • To address the semantic gap of contextual knowledge, we leverage multi-level external rationales (image, text, and combined) generated by a pretrained LVLM with chain-of-thought prompting for sarcasm detection.
  • We conduct extensive experiments and demonstrate that MiDRE outperforms existing baselines, providing both improved accuracy and interpretability in sarcasm detection.
2. Related Work
2.1. Image-Text Sarcasm Detection
Early approaches to sarcasm detection focused solely on text, leveraging lexical features (Joshi et al., 2015; Khattri et al., 2015; Joshi et al., 2016; Amir et al., 2016) and sequence modeling techniques (Zhang et al., 2016; Poria et al., 2016; Ghosh et al., 2017; Agrawal and An, 2018; Agrawal et al., 2020; Babanejad et al., 2020; Lou et al., 2021; Liu et al., 2022b). The introduction of multimodal sarcasm detection began with handcrafted image-text feature fusion (Schifanella et al., 2016). Subsequent models explored hierarchical fusion (Cai et al., 2019), cross-modal semantic reasoning (Xu et al., 2020), and attention mechanisms for capturing intra- and inter-modal incongruities (Pan et al., 2020). Graph-based methods were employed to model fine-grained token-level relations (Liang et al., 2021, 2022), while external knowledge incorporation further enhanced performance (Liu et al., 2022a). Dynamic routing strategies were later introduced for adaptive inter-modal reasoning (Tian et al., 2023). More recent work includes the MMSD2.0 dataset with refined annotations and CLIP-based modeling (Qin et al., 2023), and retrieval-augmented instruction tuning approaches (Tang et al., 2024).
While prior efforts focus on modeling multimodal incongruity and contrast, often with surface level knowledge, they often lack explicit, sarcasm-oriented reasoning signals. In this work, we leverage structured and verbose reasoning extracted via chain-of-thought (CoT) methods to guide the model to make more robust decisions via dual reasoning expert networks.
Figure 2. Rationale generation module of MiDRE. Detailed prompts are presented in appendix section A.1.
2.2. Mixture of Experts
(Jacobs et al., 1991) proposed the idea of different experts for solving sub-tasks of a larger task. In the domain of multimodal learning, several studies have used MoE approaches for improving multimodal recommendation systems (Lu et al., 2015; Ma et al., 2019; Qin et al., 2020). (Zhang et al., 2024) used MoE approach to improve the encoding performance of CLIP. (Lin et al., 2024) used soft-routers in MoE setting to achieve scaling in LLaVA model. A series of works (Wang et al., 2021; Long et al., 2023; Shen et al., 2023) assigned different experts to handle different modalities. In the domain of multimodal (audio + text + images) analysis of sarcasm, (Yu et al., 2023) used different large models as experts in audio, image or text modalities to model their interaction.
In contrast to existing approaches, we propose a novel framework that addresses the problem of sarcasm detection (image+text) from dual perspective, content and external rationales. Our dual expert approach can identify saracsm based on considering the content or external rationales or a mix of both.
3. Methodology
3.1. Task Definition
The task of image-text sarcasm detection can be framed as a binary classification problem where, given a multimodal input sample pair $x=(T,I)$, where $T$ is the textual content and $I$ is the image, the objective is to assign $x$ a label $y$ from the set $Y=\{yes,no\}$ where $yes$ means sarcastic.
3.2. Architecture of MiDRE
Our proposed model, MiDRE, detects sarcasm by reasoning from two complementary perspectives: (1) the content itself, and (2) the content enriched with external rationales. As shown in Figure 3, the internal reasoning expert ($IR$) captures inherent sarcastic cues directly from the input, using visual and textual features extracted via CLIP encoders. To identify sarcasm using both content and external knowledge, we first introduce a Rationale Generation Module that leverages a pre-trained LVLM with chain-of-thought prompting to generate structured, modality-specific sarcasm rationales across three views: image, text, and combined image-text. The external reasoning expert ($ER$) performs deeper reasoning using LLM-encoded representations of the input text and generated rationales, alongside CLIP-based image embeddings for visual context. A learnable gating network then adaptively fuses the outputs of $ER$ and the $IR$, allowing the model to balance reliance on external rationales versus direct image-text cues for more accurate and robust sarcasm detection. This combination of $IR$, $ER$ and gating is incorporated in all encoder layers of an encoder-decoder based LLM.
Figure 3. Architecture of the proposed MiDRE model.
3.3. Rationale Generation Module
This module, illustrated in Figure 2, is responsible for generating external reasoning or rationales about why a given image-text pair is sarcastic or not, by leveraging the pretrained knowledge of LVLMs. Sarcasm can emerge from either one modality alone or through an interplay between both modalities. To capture these diverse scenarios, we extend the idea of CoT prompting to the multimodal setting by sequentially guiding LVLM to reason in three independent stages: image-only, text-only, and joint image-text.
We observe that when the LVLM was prompted with both the image and text to generate all three rationales together, it often blended the sources—such as incorporating image details into the text-only rationale, and vice versa. To avoid this cross-modal interference and ensure that each rationale reflects its intended modality, we designed an independent three-stage extraction process. This setup helps maintain clear boundaries between image, text, and combined reasoning.
The three-stage process is described below.
Stage 1- Image-based rationale generation: In this stage, the LVLM is prompted to focus exclusively on the image, instructing it to describe the image content, infer possible implications, and reason whether the visual content alone could indicate sarcasm or irony. This helps capture cues like meme-style content or ironic visual contexts. This stage can be formulated as: $R^{I}=LVLM(Prompt^{I},I)$
Stage 2- Text-based rationale generation: Here, the LVLM is prompted to consider only the text to identify signs of sarcasm. This stage isolates the textual sarcasm signal without visual influence. We formulate this stage as: $R^{T}=LVLM(Prompt^{T},T)$
Stage 3- Multimodal rationale generation: In the final stage, both image and text are provided as input, and the LVLM is prompted to reason about the relationship between them. The goal of this stage is to identify cross-modal inconsistencies. We can formulate this stage as follows: $R^{M}=LVLM(Prompt^{M},I,T)$
We combine the three rationales to form a comprehensive external reasoning input $R=[R^{I},R^{T},R^{M}]$. This combined rationale encapsulates detailed external reasoning signals from each modality and their interaction. The order of combination of rationales is arbitrary as we observed very little variation in performance during our experiments.
3.4. Dual Reasoning Expert Networks
The internal reasoning expert $(IR)$, external reasoning expert $(ER)$ and the adaptive gate are deeply integrated into every encoder layers of the LLM. Once the rationale $R$ is extracted from the rationale generation module, the modified input sample becomes $x=\{I,T,\hat{T}\}$, where $\hat{T}=[T,R]$ is the combination of text and rationale. $ER$ considers the image-text along with rationales as input. The input required for $ER$ is extracted as
(1) $H_{I}=VE(I),\quad H_{\hat{T}}^{0}=E_{LLM}(\hat{T})$
where $H_{I}\in\mathbb{R}^{m\times d_{v}}$ is the image embedding from CLIP visual encoder $VE(\cdot)$, and $m$ is the number of image patches. $H_{\hat{T}}^{0}\in\mathbb{R}^{n\times d_{t}}$ is the embedding of the combined text and rationale input $\hat{T}$ from LLM embedding layer $E_{LLM}(\cdot)$, where $n$ is the token length. Since image embeddings $H_{I}$ is from the CLIP image encoder, to align it with the LLM embedding space, we use a separate learnable projection layer:
(2) $H_{I^{\prime}}=H_{I}W_{i}$
$\hat{T}$ contains richer and more complex semantic information than $T$, hence we pass it through the transformer layers of the encoder to leverage the deep contextual reasoning capabilities of the LLM and facilitate more sophisticated integration of external rationale signals.
Specifically in LLM encoder layer ${i}$, the computation of $ER$ can be formulated as:
(3) $M_{ER}^{i}=ER^{i}(H_{\hat{T}}^{i-1},H_{I^{\prime}})$
where $M_{ER}^{i}$ is the output from the external reasoning expert $ER^{i}$ in the encoder layer $i$. $ER$ is implemented like a cross-attention mechanism as follows:
(4) $ER(H_{\hat{T}},H_{I^{\prime}})=\text{softmax}\left(\frac{Q_{\hat{T}}K^{T}_{I^{\prime}}}{\sqrt{d}}\right)V_{I^{\prime}}$
where $\{Q_{\hat{T}},K^{T}_{I^{\prime}},V_{I^{\prime}}\}=\{H_{\hat{T}}W_{Q},H_{I^{\prime}}W_{K},H_{I^{\prime}}W_{V}\}$. $\{W_{Q},W_{K},W_{V}\}\in\mathbb{R}^{d\times d_{k}}$ are query, key and value projections respectively.
The internal reasoning expert ($IR$) is designed to learn semantic associations directly from the input image-text pair without relying on any external rationales. Since $T$ is comparatively simpler than the enriched rationale-augmented input $\hat{T}$, and to avoid potential interference or leakage of rationale information into the internal reasoning stream, we extract the embedding of $T$ separately using a CLIP text encoder. This design choice ensures that $IR$ operates purely on the original image and text content. To validate this design choice, we also experimented with a reversed configuration—feeding $T$ into the LLM encoder and $\hat{T}$ into the CLIP text encoder, but this led to inferior performance, as reported in Appendix A.2.
The text input required for $IR$ is extracted as $H_{T}=VE(T)$
where $H_{T}\in\mathbb{R}^{n\times d_{t}}$ is the embedding of text only input $T$ from CLIP text encoder $TE(\cdot)$. Since text embeddings $H_{T}$ is from CLIP encoder, to align it with the LLM embedding space, we use a separate learnable projection layer:
(5) $H_{T^{\prime}}=H_{T}W_{t}$
We utilize the same image embeddings ${H}_{I^{\prime}}$ as of $ER$. At each LLM encoder layer $i$, it produces the output:
(6) $M_{IR}^{i}=IR^{i}(H_{T^{\prime}},H_{I^{\prime}})$
The internal reasoning expert is realized as:
(7) $IR(H_{T^{\prime}},H_{I^{\prime}})=\text{softmax}\left(\frac{Q_{T^{\prime}}K^{T}_{I^{\prime}}}{\sqrt{d}}\right)V_{I^{\prime}}$
where $\{Q_{T^{\prime}},K_{I^{\prime}},V_{I^{\prime}}\}=\{H_{T^{\prime}}W_{Q},H_{I^{\prime}}W_{K},H_{I^{\prime}}W_{V}\}$.
While $IR$ suffices in cases where the text and image directly convey sarcastic cues, there are scenarios where IR alone is inadequate, and the model must rely on the $ER$, for contextual understanding. However, external reasoning can also be noisy or unreliable. To adaptively balance these signals, we introduce a gating mechanism that dynamically weighs the contributions of $IR$ and $ER$ at each encoder layer $i$. The gating is realized as a bottleneck layer that takes the concatenated embeddings of both expert outputs:
(8) $G^{i}=\text{softmax}(W_{g2}^{i}\cdot\text{GELU}\left(W_{g1}^{i}[M_{ER}^{i};M_{IR}^{i}]\right))$
where $W_{g1}^{i}\in\mathbb{R}^{2d_{t}\times d_{b}}$, $W_{g2}^{i}\in\mathbb{R}^{d_{b}\times 2}$, $d_{b}<d$ defines the bottleneck dimension, and $G^{i}\in\mathbb{R}^{2}$.
The final gated output is then computed as:
(9) $M_{G}^{i}=G^{i}_{1}\odot M_{ER}^{i}\;+\;G^{i}_{2}\odot M_{IR}^{i}$
The bottleneck design allows the gate to act as an information filter to dynamically attend to either $IR$, $ER$, or a context-dependent combination of both. We utilize a residual connection to generate the final output from encoder layer $i$ which will be fed to the next encoder layer $i+1$:
(10) $H_{\hat{T}}^{i}=H_{\hat{T}}^{i-1}+M_{G}^{i}$
3.5. Model Training and Prediction
The output $H_{\hat{T}}^{L}$ from the final encoder layer $L$ is fed to the LLM decoder, which generates the label from the label set $\{yes,no\}$. The LLM and the image and text projection layers are trained to minimize the cross-entropy loss between the ground truth label and the generated label. The CLIP visual and text encoders are only used as feature extractors and are frozen during training.
4. Experiments and Results
4.1. Datasets and Experimental Settings
We evaluate our proposed model MiDRE on MMSD2.0 (Qin et al., 2023) dataset. We also report the performance on MMSD(Cai et al., 2019), which is a biased dataset in Table 3. The dataset statistics are reported in Table 1. We use flan-t5-large 111google/flan-t5-large as our backbone LLM and for generating rationales, we use the LVLM Qwen2-VL-7B-Instruct [^2] Qwen/Qwen2-VL-7B-Instruct, due to its better zero-shot performance on our task over other LVLMs as validated in Figure 6. For image and text feature extractors, we use clip-vit-large-patch14333openai/clip-vit-large-patch14. We train our model for 20 epochs with Adam optimizer, maintaining a batch size of 12 and a learning rate of 5e-5. We pick the model with the best validation accuracy for testing. All experiments are carried out on 1xNvidia A100 80GB GPU.
Table 1. Statistics of datasets.
MMSD/MMSD2.0 Train Val Test
Sentences 19816/19816 2410/2410 2409/2409
Positive 8642/9572 959/1042 959/1037
Negative 1174/10230 1451/1368 1450/1372
Figure 4. Unique test samples correctly classified by MiDRE(w/o IR) vs MiDRE(w/o ER).
Table 2. Performance comparison of different models on the MMSD2.0 dataset. * denotes that we reimplemnt the models. indicates our method significantly outperforms baselines with $p<0.001$.
Model Acc. P R F1 Text-Only TextCNN (Kim, 2014) 71.61 64.62 75.22 69.52 BiLSTM (Graves and Schmidhuber, 2005) 72.48 68.02 68.08 68.05 SMSD (Xiong et al., 2019) 73.56 68.45 71.55 69.97 RoBERTa (Liu et al., 2019) 79.66 76.74 75.70 76.21 ChatGLM2-6B (Du et al., 2021) 78.41 78.15 78.65 78.23 LLaMA2-7B (Touvron et al., 2023) 82.52 82.15 82.46 82.27 Image-Only ResNet (He et al., 2015) 65.50 61.17 54.39 57.58 ViT (Dosovitskiy et al., 2020) 72.02 65.26 74.83 69.72 Multimodal HFM (Cai et al., 2019) 70.57 64.84 69.05 66.88 Att-BERT (Pan et al., 2020) 80.03 76.28 77.82 77.04 CMGCN (Liang et al., 2022) 79.83 75.82 78.01 76.90 HKE (Liu et al., 2022b) 76.50 73.48 71.07 72.25 DynRT-Net (Tian et al., 2023) 71.40 71.80 72.17 71.34 Multi-view CLIP* (Qin et al., 2023) 85.14 80.18 88.21 84.00 RAG-LLaVA* (Tang et al., 2024) 85.11 80.02 87.34 83.51 MiDRE (ours) 86.18 86.62 89.00 87.79
4.2. Baselines
We consider three groups of baselines: image-modality, text-modality and multimodal methods. For image-modality, we follow (Cai et al., 2019) and use the representations from the pooling layer of ResNet (He et al., 2015) and embedding of the [CLS] token from ViT (Dosovitskiy et al., 2020). For text-modality methods, we consider TextCNN (Kim, 2014), BiLSTM (Graves and Schmidhuber, 2005), SMSD (Xiong et al., 2019) which uses a low rank bilinear pooling using a self-matching network, RoBERTa(Liu et al., 2019) and more recent LLMs like LLaMA2-7B (Touvron et al., 2023) and ChatGLM2-6B (Du et al., 2021). For the multimodal methods, we consider HFM (Cai et al., 2019), D&R Net (Xu et al., 2020), Att-BERT (Pan et al., 2020), InCrossMGs (Liang et al., 2021), CMGCN (Liang et al., 2022), HKE (Liu et al., 2022a), DynRT-Net (Tian et al., 2023), Multi-view CLIP (Qin et al., 2023) and RAG-LLaVA (Tang et al., 2024). We do not show the results of InCrossMGs and D&R Net on MMSD2.0 due to the unavailability of their codes.
Figure 5. Examples demonstrating the importance of external rationales for sarcasm identification. ’Baseline’ indicates Multi-view CLIP (the best performing baseline for MMSD2.0), ER(IR) Gate Values represent the importance given to external (internal) expert module by the gating network in MiDRE.
4.3. Main Results
Following (Cai et al., 2019), we use Accuracy (Acc.), macro-average precision (P), macro-average recall (R), and macro-average F1 (F1) to report the performance of our model in Table 5. Our main observations are follows: (1) Our method MiDRE outperforms all SOTA baselines in MMSD2.0: On MMSD2.0 dataset, MiDRE outperforms all baseline across unimodal and multimodal settings. Specifically it ourpeforms the best SOTA multimodal baseline Multi-view CLIP by good margins across metrics (Acc: $\uparrow 1.04\%$, P: $\uparrow 6.44\%$, R: $\uparrow 0.79\%$, F1: $\uparrow 3.79\%$ ). This proves the efficacy of using structured rationales and having dual reasoning paths to identify sarcasm.
(2) MiDRE outperforms all multimodal baselines that use external knowledge: We notice that our method significantly outperform baselines like HFM (Cai et al., 2019) (Acc: $\uparrow 15.61\%$ in MMSD2.0) which uses image attributes generated using ResNet, CMGCN (Liang et al., 2022) (Acc: $\uparrow 6.35\%$ in MMSD2.0) which utilizes object-attribute pairs from a detectron module and HKE (Liu et al., 2022a) (Acc: $\uparrow 9.68\%$ in MMSD2.0) which incorporates image caption, image attributes and adjective-noun pairs from image. This highlights the importance of leveraging structured, verbose and sarcasm-oriented knowledge.
4.4. Results on Biased MMSD Dataset
MMSD is a text-biased dataset, as demonstrated by Qin et al. (2023). This bias arises from the presence of strong textual cues, such as hashtags and emojis that predominantly occur in one class. Since text-only models do not process visual information, they can directly exploit these cues without interference from the image modality, often outperforming multimodal models on this dataset. For completeness, we evaluate MiDRE on MMSD. As shown in Table 3, the text-only LLM ChatGLM2-6B even outperforms existing multimodal models. Notably, traditional language models like RoBERTa perform on par with larger LLMs such as LLaMA2-7B and ChatGLM2-6B, further emphasizing the dataset’s inherent bias. Despite this, MiDRE achieves consistently better performance than all prior multimodal approaches by significant margins, demonstrating its robust reasoning ability. Moreover, the small performance gap between MiDRE and ChatGLM2-6B suggests that MiDRE effectively leverages strong textual signals while also integrating complementary visual and external knowledge.
5. Comparison with LVLMs
We compare the performance of state-of-the-art large vision-language models (LVLMs) under a zero-shot prompting setup for MMSD2.0, as reported in Figure 6. Despite advanced zero-shot capabilities and contextual knowledge, these models perform poorly on our task. Since fine-tuning of such large models is expensive, we press upon the need of developing smaller models like MiDRE that are explicitly designed for nuanced reasoning tasks such as multimodal sarcasm detection. At the same time, it reveals how the capabilities of such smaller models can be significantly amplified by strategically leveraging contextual understanding from large vision-language models.
6. Ablation
To further prove the efficacy of our approach, we answer the following research questions: (1) What is the impact of having only one expert network? (2) What advantage does the gating network provide? (3) Do we need rationales from image, text and multimodal views?
6.1. Single Expert Network
To better understand the contributions of the internal (IR) and external (ER) reasoning experts, we perform an ablation study on MiDRE by selectively removing each component (Table 4, rows 3 & 4). Removing the ER module (w/o ER) leads to a significant drop in accuracy (↓4.72% on MMSD, ↓3.91% on MMSD2.0), highlighting the importance of external contextual rationales that are not directly inferable from the input. Conversely, removing IR (w/o IR) also reduces performance (↓2.75% on MMSD, ↓2.12% on MMSD2.0), showing that ER alone struggles to compensate for noisy or incomplete rationales.
Interestingly, from Figure 4, we can observe that w/o ER correctly classifies 111 unique samples and w/o IR classifies 154 different ones, suggesting that both experts offer complementary and independently valuable reasoning capabilities. Moreover, the larger performance drop without ER indicates its dominant role, but IR remains essential for grounding predictions in the raw image-text content, enabling more reliable sarcasm detection through their combined use.
Table 3. Performance comparison of different models on the MMSD dataset. * denotes that we reimplement the models.
Model Acc. P R F1 Text-Only TextCNN (Kim, 2014) 80.03 74.29 76.39 75.32 BiLSTM (Graves and Schmidhuber, 2005) 81.90 76.66 78.42 77.53 SMSD (Xiong et al., 2019) 80.90 76.46 75.18 75.82 RoBERTa (Liu et al., 2019) 93.97 90.39 94.59 92.45 ChatGLM2-6B (Du et al., 2021) 94.02 93.46 94.14 93.76 LLaMA2-7B (Touvron et al., 2023) 93.97 93.42 94.09 93.72 Image-Only ResNet (He et al., 2015) 64.76 54.41 70.80 61.53 ViT (Dosovitskiy et al., 2020) 67.83 57.93 70.07 63.40 Multimodal HFM (Cai et al., 2019) 83.44 76.57 84.15 80.18 D&R Net (Xu et al., 2020) 84.02 77.97 83.42 80.60 Att-BERT (Pan et al., 2020) 86.05 80.87 85.08 82.92 InCrossMGs (Liang et al., 2021) 86.10 81.38 84.36 82.84 CMGCN (Liang et al., 2022) 86.54 - - 82.73 HKE (Liu et al., 2022b) 87.36 81.84 86.48 84.09 DynRT-Net (Tian et al., 2023) 93.59 93.06 93.60 93.31 Multi-view CLIP* (Qin et al., 2023) 88.33 82.66 88.65 88.55 RAG-LLaVA* (Qin et al., 2023) 89.97 89.26 89.58 89.42 MiDRE (ours) 93.81 93.11 93.84 93.47
Figure 6. Zero-shot Performance of LVLMs on MMSD2.0 test samples.
Figure 7. Examples demonstrating how noisy rationales can provide contextual cues. ’Baseline’ indicates Multi-view CLIP (the best performing baseline for MMSD2.0), ER(IR) Gate Values represent the importance given to external (internal) expert module by the gating network in MiDRE.
6.2. Impact of Gating Network
To understand the impact of the gating network, we ablate by removing it from MiDRE, reported in Table 4 as w/o Gate. In this case, we simply add the output from the internal reasoning expert and external reasoning expert. We notice that performance drops notably (Acc: $\downarrow 3.26\%$ in MMSD and $\downarrow 1.34\%$ in MMSD2.0). This decrease in performance indicates that the gating network plays a crucial role in controlling the interaction between the internal reasoning network (IR) and external reasoning module (ER). The adaptive gating ensures that the model relies on the most informative reasoning path to make a decision. In cases where it can’t rely on one single reasoning path, it happens to look at both reasoning paths to arrive at a decision, which is described in detail in Section 7. We also validate our design choice of using a bottleneck architecture for the gating mechanism by comparing it against a simpler linear gate, as reported in Table 4 (w Linear Gate). The bottleneck gate consistently performs better, suggesting that it facilitates more expressive interactions between the expert outputs.
6.3. Need of All Three Rationales
To validate the necessity of all three rationales (text, image and multimodal) in MiDRE, we conduct a systematic ablation study presented in Table 4 (rows 1 & 2). When only the multimodal rationale is used, the model’s performance drops significantly,(Acc: $\downarrow 3.68\%$ in MMSD and $\downarrow 2.12\%$ in MMSD2.0), indicating that relying solely on the multimodal rationale is insufficient. Similarly, when only the text and image rationale is used, the performance drops (Acc: $\downarrow 3.14\%$ in MMSD and $\downarrow 1.5\%$ in MMSD2.0), compared to the full model. In contrast, the full MiDRE model, which combines the image, text and multimodal rationales achieves the highest performance , emphasizing the necessity of views from both image and text modalities as well as the view from the interaction between them.
7. Qualitative Analysis
We qualitatively explore the following research questions: (1) How does external rationale help? (2) How does noisy external rationale affect model performance?
Table 4. Ablation of our model MiDRE. w/o means without that specific component.
MiDRE MMSD MMSD2.0 Acc. P R F1 Acc. P R F1 Only Mul. Rationale 90.13 87.35 90.41 88.85 84.06 80.51 86.69 83.48 Only Txt+Img Rationale 90.67 89.21 88.37 88.78 84.68 82.11 82.35 82.23 w/o ER 89.09 85.82 89.11 87.43 82.27 74.51 89.39 81.28 w/o IR 91.06 89.62 88.45 89.03 84.06 80.03 83.89 81.92 w/o Gate 90.55 85.13 89.97 87.48 84.84 78.96 88.33 83.38 w Linear Gate 91.11 85.82 90.02 87.86 84.89 79.17 88.52 83.58 MiDRE 93.81 93.11 93.84 93.47 86.18 86.62 89.00 87.79
7.1. Need of External Rationale
External reasoning is crucial in sarcasm detection, particularly when textual and visual cues alone fail to establish contradiction. For instance, in Figure 5(a), the image shows Easter-themed products on a shelf, while the text questions when Halloween products will arrive, implying that Easter items are being sold too early. A model without external knowledge may struggle to detect sarcasm since the image-text itself does not explicitly indicate any contradiction. However, external reasoning clarifies that Easter eggs appearing in January is unusual, reinforcing the sarcasm. MiDRE successfully classifies this instance as sarcastic, with a high gate value for External Rationale (ER) module (0.876), while the baseline and MiDRE without ER fail. In Figure 5(b), the text references a famous sports trade (Mike Ditka’s excessive trade for Ricky Williams) to suggest that the Broncos should do something similarly extreme. The accompanying image with confused expressions enhances the sarcasm. Without knowledge of the Ditka trade, the sarcasm might be missed, as the text could be interpreted as a genuine suggestion. External reasoning bridges this gap by explaining the historical context, leading MiDRE to detect sarcasm correctly (ER gate value: 0.921). In the case of Figure 5(c), both the image and text lack strong sarcastic elements. The image depicts a casual mirror selfie, and the text states that the person probably takes selfies with filters. MiDRE assigns a high ER gate value (0.781), suggesting that the Internal Reasoning (IR) module might be uncertain due to the overtly positive tone of the text. However, ER helps by recognizing that there is no cultural or contextual contradiction that would indicate sarcasm.
7.2. Impact of Noisy Rationale
External rationales, even when noisy or indirect, enhance the internal reasoning (IR) module by providing valuable contextual cues for sarcasm detection. For example, in Figure 7(a), MiDRE without external reasoning (ER) fails to detect sarcasm, while MiDRE succeeds by leveraging the image rationale, which highlights the neutral design of the seat. This neutrality amplifies the exaggerated praise in the text, making the sarcasm more apparent. Similarly, in 7(b), although the image rationale doesn’t directly contradict the caption, it describes the dog’s playful behavior, subtly clashing with the claim of it being ”well-behaved” at the vet. In both cases, ER offers factual descriptions that help IR identify subtle incongruities, demonstrating the complementary role of external rationales in improving sarcastic intent recognition.
Figure 8. Examples of misclassified samples
8. Error Analysis
Our model, MiDRE, fails in cases where both internal and external reasoning fall short. In Figure 8(a), the image explicitly mentions ”sarcasm,” leading to model bias—falsely associating the term with sarcasm without deeper analysis. In Figure 8(b), sarcasm emerges from the contradiction of the makeshift arrangement of sticking a calendar to the wall using a tape with the caption, “I’m really classy.” Here, internal reasoning struggles to detect the subtle contradiction, while external reasoning lacks the context to flag it, causing the model to miss the sarcasm.
9. Conclusion
This work addresses the limitations of existing approaches that use shallow external knowledge, which is often surface-level and generic in nature. We propose MiDRE, a novel framework that combines internal reasoning over image-text pairs with external reasoning guided by structured rationales generated through CoT prompting to a LVLM. MiDRE leverages a learnable gating mechanism to dynamically integrate both reasoning pathways, allowing the model to adaptively rely on commonsense and contextual knowledge when necessary. Extensive experiments on multiple benchmarks demonstrate that MiDRE achieves state-of-the-art performance and provides more interpretable and robust sarcasm detection, especially in complex scenarios that require understanding beyond literal visual or textual cues. This highlights the importance of structured external rationales and expert-guided reasoning in advancing multimodal understanding tasks like sarcasm detection.
10. Appendix
10.1. Detailed Prompts
This section outlines the prompts in full details that we use for the rationale generation module.
10.1.1. Image Prompt
You are given an image. Follow the steps below:### Task 1: Image Analysis- Carefully analyze the image, considering any visible elements, context, and any potential irony or contrast in the visual content.- Based on your analysis, determine if the image contains sarcasm.  - Print **1** if the image is sarcastic.  - Print **0** if the image is not sarcastic.### Task 2: Reasoning- Provide a **concise** reasoning for why you chose the label in Task 1.- Ensure that your reasoning explains any ironic or sarcastic elements or the lack thereof in the image.### Output Format:- **Label**: [1/0]- **Reason**: [<reason>]### Notes:- The **Label** is determined by whether you believe the image is sarcastic or not.- The **Reason** should be a **concise** explanation directly related to the content of the image.- The tasks should be handled **independently**, but logically connected (task 2 explains task 1).- Your output should **strictly follow** this format without any additional commentary or explanation:# Adhere strictly to the order and independence of tasks as described. Provide outputs exactly in the specified format.
10.1.2. Text Prompt
You are given a text. Follow the steps below:### Task 1: Image Analysis- Carefully analyze the text, considering any exaggeration, incongruity or ironic elements in the textual content.- Based on your analysis, determine if the text contains sarcasm.  - Print **1** if the text is sarcastic.  - Print **0** if the text is not sarcastic.### Task 2: Reasoning- Provide a **concise** reasoning for why you chose the label in Task 1.- Ensure that your reasoning explains any ironic or sarcastic elements or the lack thereof in the text.### Output Format:- **Label**: [1/0]- **Reason**: [<reason>]### Notes:- The **Label** is determined by whether you believe the text is sarcastic or not.- The **Reason** should be a **concise** explanation directly related to the content of the text.- The tasks should be handled **independently**, but logically connected (task 2 explains task 1).- Your output should **strictly follow** this format without any additional commentary or explanation:# Adhere strictly to the order and independence of tasks as described. Provide outputs exactly in the specified format.
10.1.3. Multimodal Prompt
You are provided with an **Image** and a **Supporting Text**. Perform the following tasks independently:# **Definitions**:# - **Supporting Text**: The text provided separately, not embedded in the image.# - **Caption**: The text embedded within the image.1. **Extract Caption**:   - Extract the caption (if any) from the image. This step should not involve analyzing sarcasm.2. **Analyze Supporting Text**:   - Consider only the provided supporting text and ignore the caption or any information from the image.   - Is the supporting text sarcastic? Answer `1` for sarcastic and `0` for non-sarcastic.   - Provide reasoning based only on the supporting text.3. **Analyze Image and Caption**:   - Analyze the image and its caption together, considering any contrasts, contradictions, or ironic elements.   - Is the image-caption pair sarcastic? Answer `1` for sarcastic and `0` for non-sarcastic.   - Provide reasoning based on the image-caption pair.**Output Format**:Caption in Image: [Caption]Supporting Text Analysis:  - Label: [1/0]  - Reason: [<reason>]Image-Caption Analysis:  - Label: [1/0]  - Reason: [<reason>]**Guidelines**:- Keep the analysis of the supporting text completely independent of the image and caption.- Provide outputs in the specified format without additional commentary.- Clearly separate reasoning for the supporting text and the image-caption pair.
10.2. Performance of Reverse Configuration
Table 5. Performance on the MMSD2.0 dataset.
Model Acc. P R F1 MiDRE (w reverse) 82.13 81.62 80.20 80.90